OpenStudy (anonymous):

The line of symmetry of the parabola whose equation is y = ax2 - 4x + 3 is x = -2. What is the value of "a"? A.) -2 B.) -1 or C.) -1/2

OpenStudy (anonymous):

y=a(−2)^2−4(−2)+3 Squaring an expression is the same as multiplying the expression by itself 2 times. y=(a⋅(−2)(−2))−4(−2)+3 Multiply to simplify the expression (−2)(−2). y=(a⋅4)−4(−2)+3 Multiply a by 4 to get 4a. y=(4a)−4(−2)+3 Multiply −4 by each term inside the parentheses. y=(4a)+8+3 Remove the parentheses around the expression 4a. y=4a+8+3 Add 3 to 8 to get 11. y=4a+11 Since a is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. 4a+11=y Since 11 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 11 from both sides. 4a=−11+y Move all terms not containing a to the right-hand side of the equation. 4a=y−11 Divide each term in the equation by 4. a=14y−114

OpenStudy (anonymous):

Consider finding the roots of \[ax ^{2} - 4x +3 = 0\] With a=-2 they are \[-1 +- \frac{ \sqrt{10} }{ 2 }\] With a = -1 they are \[-2+- \sqrt{7}\] With a = -1/2 they are \[-4 +- \sqrt{22}\] In the second case, the roots are symmetric about x=-2; therefore, the parabola is too.

OpenStudy (anonymous):

Thank You BillBell!

OpenStudy (anonymous):

Welcome!

OpenStudy (anonymous):

... and thank you for the medal.